# The learned range test method for the inverse inclusion problem

### v1

@@ 1. Introduction
Over the past decade, deep learning methods have attracted extensive attention because of their salient success in solving inverse problems; see, for example, [23, 61, 59, 54, 15]. Directly solving inverse problems by constructing an end-to-end neural network and ignoring the physical model loses interpretability and often leads to less effective reconstruction methods [9, 47]. As a result, the integration of conventional methods with deep learning methods has become a popular direction in recent years, also in the context of qualitative methods. For example, in [30], the authors propose a deep learning method for electrical impedance tomography based on the conventional direct sampling method. They employ fully connected neural networks and convolutional neural networks to approximate the indicator function of the inclusion. The study [43] integrates the orthogonality sampling method with deep neural networks to reconstruct the geometry of a penetrable object from scattering measurements generated by a single incident wave. In [50], the authors propose a sampling-based deep learning method to tackle the inverse medium scattering problem. Specifically, they integrate the results generated by the direct sampling method with a U-Net to learn the relationship between the indicator function and the true resolution of the target. In [45], the authors consider the inverse inhomogeneous medium problem using a learning-based iterative algorithm. Briefly speaking, they make use of a deep neural network to learn the a priori information on the shape of the unknown scatterer, and then refine the reconstruction by using the iteratively regularized Gauss-Newton method.
@@
The remainder of this paper is structured as follows. The mathematical formulation of our inverse problem is stated in Section 2, where a literature review concerning this problem is also included. In Section 3, we briefly revisit the RT from the theoretical and numerical perspectives. Section 4 is devoted to the LRT method, where the architecture of the neural network and the three-step strategy are described. Various numerical experiments are carried out in Section 5. In Section 6, we provide some concluding remarks. Finally, some details on the computational aspects, with a particular focus on the training of the neural networks of this work, are discussed in Appendix A.
@@ 2.2. Literature review
In general, iterative methods yield relatively good reconstructions via the iterative procedure. However, their drawbacks are the heavy computational burden and the need of many a priori information about the target, which might not be available in practice. Therefore, as a remedy, qualitative methods were introduced in order to obtain some limited information of the target. Some popular qualitative methods include the sampling method [44, 22], the enclosure method [35, 36], the probe method [34] and the MUSIC-type method [7]. We also refer the reader to [39, 19, 40, 52] for additional qualitative methods and the relative technical details. Compared with the iterative methods, the qualitative methods allow for a good trade-off between accuracy and computational cost.
@@ 5.1. Setup
In all the numerical experiments, we restrict ourselves to the case where $n=2$, with the background medium $\Omega$ being a disk centered at the origin with radius $R_{\Omega}=10$. The initial searching area $\tilde{\Omega}=[-2,2]\times[-2,2]$ is divided into a uniform $41\times 41$ Cartesian grid $\Gamma$, so that $N=41^{2}$. This is the same setup used in [56], see also Figure 2, which allows us to better compare the results. For every inclusion, the synthetic data $\partial_{\nu}\omega$ is obtained by solving the forward problems (1) and (2) with boundary value $g(x,y)=xy$, $(x,y)\in\partial\Omega$. We use $\partial_{\nu}\omega$ at $\mathcal{N}=400$ collocation points on $\partial\Omega$ as the measurement. To demonstrate the robustness of our strategy, we consider the following noisy data with the noise level $\delta$:
@@ 5.2. Construction of the dataset
$h+r(\cos\psi_{j},\sin\psi_{j}),\qquad j=1,\dots,k.$
This construction guarantees that the inclusions are convex polygons satisfying the distance property (5). The test sets are generated in a similar way.
@@ 5.4. Results
which measures the distance between the actual contrast $\chi(B)$ and the predicted result $\Phi_{\Theta^{*}}(y^{\delta}(B))$ of the LRT, for a domain $B$ in the test set. In all the figures, the red polygons or curves refer to our reconstruction targets, and the corresponding MSE is shown below each picture.
(a) 0.0059
(b) 0.0089
(c) 0.0071
(d) 0.0111
(e) 0.0158
(f) 0.0073
(g) 0.0060
(h) 0.0092
(i) 0.0093
(j) 0.0059
(k) 0.0120
(l) 0.0089
Figure 6. Some samples of the reconstructions of Scenario 1. The mean squared error (13) is recorded below each image.
@@
  • Scenario 3: Evaluate the generalization performance of the LRT. We consider $4$ inclusions that are not within the class of our training set, in order to test the generalization performance of the LRT. These four inclusions are used as the test samples in Scenario 1 and Scenario 2. The inclusions and the reconstructions are depicted in Figure 9.
From Figures 6 and 8, we can observe that the LRT can reconstruct the inclusion with high accuracy in both the noise-free and noisy measurement cases. In particular, it recovers the shape and the position of the inclusion even when the target is small and far away from the boundary; see, for instance, Figure 6(c) and Figure 8(c). As shown in Figures 7(a) and 7(b), the loss values for the samples in the test set are primarily distributed within the interval $[0,0.02]$, which further illustrates the effectiveness of the LRT. Furthermore, the numerical results shown in Figures 9 suggest that the LRT exhibits excellent generalization property. It is noteworthy that only one boundary measurement is employed to realize the reconstruction.
@@
(a) 0.0077
(b) 0.0155
(c) 0.0077
(d) 0.0120
(e) 0.0160
(f) 0.0054
(g) 0.0095
(h) 0.0149
(i) 0.0224
(j) 0.0089
(k) 0.0232
(l) 0.0202
Figure 8. Some samples of the reconstructions of Scenario 2. The mean squared error (13) is recorded below each image.
(a) 0.0225
(b) 0.0105
(c) 0.0071
(d) 0.0303
(e) 0.0234
(f) 0.0101
@@ 5.5.2. End-to-end fully connected neural network
It is natural to compare the LRT with a standard end-to-end fully connected neural network. For a more faithful comparison, we use the same three-step strategy of the LRT, namely, we design $L=3$ fully connected NNs with four layers, where the number of the neurons are $400$, $1200$, $1200$ and $1681$, respectively. The activation functions between the four layers are the ReLUs. The NNs take as input the measurements $y^{\delta}(B)$ and output an approximation of $\chi(B)$, for $B\in\mathcal{T}_{\ell}$. We use the same reconstruction algorithm discussed in §$\S$4.3, in which the learned classifier is combined with the learned fully connected NNs (which replace the $\Phi_{\theta_{\ell}}$). The numerical results of the same inclusions considered in the previous case (§$\S$5.5.1) are shown in Figures 10 and 11, third row. We also test other fully connected NNs with different depths and widths, but they turn out to give similar results. As it is evident, the results are much worse than those of the LRT, and the NNs tend to favor the value 00 in the reconstructions. This indicates that the architecture of the networks used in the LRT, which takes into account the physical model and the RT method, is a key aspect for a successful recovery.
@@ 6. Conclusions
The method discussed in this work is adapted to one single measurement and one single inclusion to detect. It would be natural to extend this approach to more complicated disconnected inclusions, perhaps by using multiple boundary measurements. Furthermore, it would be interesting to investigate whether the method discussed in this paper, based on a “write as a NN and then learn” approach, can be extended to other qualitative methods for inverse problems. Finally, more extensive numerical results, possibly in real-world scenarios, are needed to fully evaluate the capabilities of this technique. These investigations are left for future research.
@@ Acknowledgements
This work has been carried out at the Machine Learning Genoa (MaLGa) center, University of Genoa. S.S. gratefully acknowledges the support provided by the China Scholarship Council program (Project ID: 202306090232), the National Natural Science Foundation of China (Nos. 12471395, 12071072) and the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant No. BK20233002. G.S.A. is supported by the Air Force Office of Scientific Research under award number FA8655-23-1-7083 and by the European Union (ERC, SAMPDE, 101041040). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. G.S.A. is a member of the “Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni”, of the “Istituto Nazionale di Alta Matematica”. The research was supported in part by the MIUR Excellence Department Project awarded to Dipartimento di Matematica, Università di Genova, CUP D33C23001110001. Co-funded by European Union – Next Generation EU.

### v2

@@ 1. Introduction
Over the past decade, deep learning methods have attracted extensive attention because of their salient success in solving inverse problems; see, for example, [26, 70, 66, 60, 17]. Directly solving inverse problems by constructing an end-to-end neural network and ignoring the physical model loses interpretability and often leads to less effective reconstruction methods [9, 52]. As a result, the integration of conventional methods with deep learning methods has become a popular direction in recent years, also in the context of qualitative methods. For example, in [34], the authors propose a deep learning method for electrical impedance tomography based on the direct sampling method. They employ fully connected neural networks and convolutional neural networks to approximate the indicator function of the inclusion. The study [47] integrates the orthogonality sampling method with deep neural networks to reconstruct the geometry of a penetrable object from scattering measurements generated by a single incident wave. In [55], the authors propose a sampling-based deep learning method to tackle the inverse medium scattering problem. Specifically, they integrate the results generated by the direct sampling method with a U-Net to learn the relationship between the indicator function and the true resolution of the target. In [49], the authors consider the inverse inhomogeneous medium problem using a learning-based iterative algorithm. Briefly speaking, they make use of a deep neural network to learn the a priori information on the shape of the unknown scatterer, and then refine the reconstruction by using the iteratively regularized Gauss-Newton method. Inspired by the linear sampling method in [10], the work [51] employs kernel methods for inverse source and scattering problems, which are modeled by integral operators with specific kernels. Inclusion reconstruction may also be viewed as an imaging problem, for which convolutional neural networks (CNN) are often used because of their capability to capture spatial structures. Several CNN-based methods have been proposed to solve this kind of inverse problems, see e.g. [24, 69, 56].
@@
The remainder of this paper is structured as follows. The mathematical formulation of our inverse problem is stated in Section 2, where a literature review concerning this problem is also included. In Section 3, we briefly revisit the RT from the theoretical and numerical perspectives. Section 4 is devoted to the LRT method, where the architecture of the neural network and the three-step strategy are described. Various numerical experiments are carried out in Section 5. In Section 6, we provide some concluding remarks. Finally, some details on the computational aspects, with a particular focus on the training of the neural networks of this work, and partial numerical results are discussed in the appendix.
@@ 2.2. Literature review
In general, iterative methods yield relatively good reconstructions via the iterative procedure. However, their drawbacks are the heavy computational burden and the need for extensive a priori information about the target, which might not be available in practice. Therefore, as a remedy, qualitative methods were introduced in order to obtain some limited information of the target. Some popular qualitative methods include the sampling method [48, 25], the enclosure method [39, 40], the probe method [38] and the MUSIC-type method [7]. We also refer the reader to [43, 21, 44, 58] for additional qualitative methods and the relative technical details. Compared with the iterative methods, the qualitative methods allow for a good trade-off between accuracy and computational cost.
@@ 5.1. Setup
In all the numerical experiments, we restrict ourselves to the case where $d=2$, with the background medium $\Omega$ being a disk centered at the origin with radius $R_{\Omega}=10$. The initial searching area $\tilde{\Omega}=[-2,2]\times[-2,2]$ is divided into a uniform $41\times 41$ Cartesian grid $\Gamma$, so that $N=41^{2}$. This is the same setup used in [63], see also Figure 2, which allows us to better compare the results. For every inclusion, the synthetic data $\partial_{\nu}\omega$ is obtained by solving the forward problems (1) and (2) with boundary value $g(x,y)=xy$, $(x,y)\in\partial\Omega$. We use $\partial_{\nu}\omega$ at $\mathcal{N}=400$ collocation points on $\partial\Omega$ as the measurement. To demonstrate the robustness of our strategy, we consider the following noisy data with the noise level $\delta$:
@@ 5.2. Construction of the dataset
(11) $h+r(\cos\psi_{j},\sin\psi_{j}),\qquad j=1,\dots,k.$
This construction guarantees that the inclusions are convex polygons satisfying the distance property (5). The test sets are generated in a similar way. The only difference between the two procedures is the strategy used to select the parameter $h$. Specifically, to generate the test set, $h$ is sampled uniformly at random in the interval $[r_{0},r_{3})$. The purpose of the classification step is precisely to classify the inclusions depending on their distance from the boundary.
@@ 5.4. Results
which measures the distance between the actual contrast $\chi(B)$ and the predicted result $\Phi_{\Theta^{*}}(y^{\delta}(B))$ of the LRT, for a domain $B$ in the test set. To further quantify the reconstruction quality, we employ the Structural Similarity Index Measure (SSIM)—an established metric used to gauge the similarity of two images [68]. This dimensionless quantity exhibits a theoretical upper bound of $1$ ($\text{SSIM}\in[-1,1]$) and demonstrates superior correlation to human visual perception, where values approaching $1$ indicate higher structural fidelity between reconstructed and ground-truth inclusions. In all the figures, the red polygons or curves refer to our reconstruction targets, and the corresponding MSE (left) and SSIM (right) are shown below each picture.
(a) 0.0059, $0.9345$
(b) 0.0089, $0.9070$
(c) 0.0071, $0.9218$
(d) 0.0111, $0.8862$
(e) 0.0158, $0.8640$
(f) 0.0073, $0.9223$
(g) 0.0060, $0.9320$
(h) 0.0092, $0.9032$
(i) 0.0093, $0.9160$
(j) 0.0059, $0.9241$
(k) 0.0120, $0.8984$
(l) 0.0089, $0.9098$
Figure 6. Some samples of the reconstructions of Scenario 1. The MSE (left) and SSIM (right) are recorded below each image.
@@
  • Scenario 3: Evaluate the generalization performance of the LRT. We consider $3$ inclusions that are not within the class of our training set. More specifically, the samples in the training set are all convex polygons (with the maximum number of sides being $8$), while these three inclusions either are non-convex or have $C^{2}$ boundaries. Notably, the diameter of the peanut-shaped inclusion is $1.5$, exceeding the maximum diameter of all the samples in the training set, which is 1.2 according to (11). To test the generalization performance of the LRT, these three inclusions are used as the test samples in Scenario 1 and Scenario 2. Here, we highlight that the selected samples are mainly used to test the generalization of LRT in terms of different shapes instead of positions. The inclusions and the reconstructions are depicted in Figure 9.
From Figures 6 and 8, we can observe that the LRT can reconstruct the inclusion with high accuracy in both the noise-free and noisy measurement cases. In particular, it recovers the shape and the position of the inclusion even when the target is small and far away from the boundary; see, for instance, Figure 6(c) and Figure 8(c). It is worth recalling that the recovery of inclusions from boundary measurements is a severely ill-posed problem, for which only logarithmic-type stability holds [61, 4, 28]. As a consequence, the reconstructions obtained typically have very low resolution, unless specific methods designed for polygonal inclusions are used [12]. Thus, in general, reconstructing the exact shape of the inclusion is very difficult, and the reconstructions shown in Figures 6 and 8 are to be considered very accurate in this context.
As shown in Figures 7(a) and 7(b), the loss values for the samples in the test set are primarily distributed within the interval $[0,0.02]$, which further illustrates the effectiveness of the LRT. Furthermore, the numerical results shown in Figures 9 suggest that the LRT exhibits excellent generalization property. It is noteworthy that only one boundary measurement is employed to realize the reconstruction.
@@
(a) 0.0077, $0.9168$
(b) 0.0155, $0.8537$
(c) 0.0077, $0.8956$
(d) 0.0120, $0.8922$
(e) 0.0160, $0.8485$
(f) 0.0054, $0.9365$
(g) 0.0095, $0.9037$
(h) 0.0149, $0.8444$
(i) 0.0224, $0.7857$
(j) 0.0089, $0.9021$
(k) 0.0232, $0.7851$
(l) 0.0202, $0.7926$
Figure 8. Some samples of the reconstructions of Scenario 2. The MSE (left) and SSIM (right) are recorded below each image.
(a) 0.0225, $0.7671$
(b) 0.0105, $0.8850$
(c) 0.0071, $0.9232$
(d) 0.0303, $0.8091$
(e) 0.0234, $0.8015$
(f) 0.0101, $0.9343$
@@ 5.5.2. End-to-end fully connected neural network
It is natural to compare the LRT with a standard end-to-end fully connected neural network. For a more faithful comparison, we use the same three-step strategy of the LRT, namely, we design $L=3$ fully connected NNs with four layers, where the number of the neurons are $400$, $1200$, $1200$ and $1681$, respectively. The activation functions between the four layers are the ReLUs. The NNs take as input the measurements $y^{\delta}(B)$ and output an approximation of $\chi(B)$, for $B\in\mathcal{T}_{\ell}$. We use the same reconstruction algorithm discussed in §$\S$4.3, in which the learned classifier is combined with the learned fully connected NNs (which replace the $\Phi_{\theta_{\ell}}$). The numerical results of the same inclusions considered in the previous case (§$\S$5.5.1) are shown in Figures 10 and 11, third row. We also test other fully connected NNs with different depths and widths, as well as CNNs because of their relevance for imaging tasks, but they turn out to give similar results, which we omit for the sake of conciseness. As it is evident, the results are much worse than those of the LRT, and the NNs tend to favor the value 00 in the reconstructions. This indicates that the architecture of the networks used in the LRT, which takes into account the physical model and the RT method, is a key aspect for a successful recovery.
@@ 6. Conclusions
Let us discuss some of the limitations of the current work. The method discussed in this paper is adapted to one single measurement and one single inclusion to detect. It would be natural to extend this approach to more complicated disconnected inclusions, perhaps by using multiple boundary measurements. Similarly, our method is primarily designed for the reconstruction of convex polygonal inclusions, based on the properties of the RT. When applied to inclusions with more complex geometries, the current approach may not achieve high accuracy as it does now. However, it can serve as an effective initial guess for iteration-based methods to further refine the reconstruction. Furthermore, the proposed strategy currently relies on training two separate neural networks, which increases the complexity of the implementation. A promising direction for improvement would be to integrate the two networks into a unified architecture, potentially through the design of a more sophisticated network structure.
It would be interesting to investigate whether the method discussed in this paper, based on a “write as a NN and then learn” approach, can be extended to other qualitative methods for inverse problems. Finally, more extensive numerical results, possibly in real-world scenarios, are needed to fully evaluate the capabilities of this technique. These investigations are left for future research.
@@ Acknowledgements
This work has been carried out at the Machine Learning Genoa (MaLGa) center, University of Genoa. S.S. gratefully acknowledges the support provided by the China Scholarship Council program (Project ID: 202306090232), the National Natural Science Foundation of China (Nos. 12471395, 12071072) and the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant No. BK20233002. G.S.A. is supported by the Air Force Office of Scientific Research under award number FA8655-23-1-7083 and by the European Union (ERC, SAMPDE, 101041040). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. G.S.A. is a member of the “Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni”, of the “Istituto Nazionale di Alta Matematica”. The research was supported in part by the MIUR Excellence Department Project awarded to Dipartimento di Matematica, Università di Genova, CUP D33C23001110001. Co-funded by European Union – Next Generation EU, Missione 4 Componente 1 CUP D53D23005770006 and CUP D53D23016180001.
@@ A.3. Computational cost
Appendix B Histograms of SSIM distribution
In Figure 13, we show the histograms of the SSIM distribution for the reconstructions obtained by the LRT, the LRT with random initialization and the RT with without learning. The results obtained are similar to those shown in Figures 7 and 12 for the MSE.
(a) LRT, $\delta=0$.
(b) LRT, $\delta=3\%$.
(c) LRT (Random initialization), $\delta=0$.
(d) LRT (Random initialization), $\delta=3\%$.
(e) Deterministic RT, $\delta=0$.
(f) Deterministic RT, $\delta=3\%$.
Figure 13. The SSIM distribution for all the samples in the test set with the reconstruction obtained with different methods. First row: the LRT. Second row: the LRT with random initialization during training. Third row: the RT (deterministic, no learning).